A saturated solution of silver benzoate AgOCOC6H5 has a pH of 8.67. The Ka for benzoic acid is 6.5x10-5. What is the Ksp for silver benzoate based on this data?
1
Expert's answer
2018-04-03T09:07:20-0400
pOH = 14-8.67 = 5.33

[OH-] = 10-pOH = 10-5.33 = 4.677·10-6

C6H5COO- + H2O → C6H5COOH + OH-

Ka = [4.677·10-6]· [4.677·10-6]/[x]· [4.677·10-6]

X = 7.2·10-2

Ksp = x2 = 5.2·10-3